I need help doing this problem:
(4x + 3w^2)(4x - 3w^2)

Answer by tmarienatree(17)

you would FOIL (First Outer Inner Last) that.
so if you look at the promblem it would go something like this!
(4x+3w^2) (4x-3w^2)
First- 4x times 4x =16x you never add or subtract the exponets!
Outer- 4x times 3w^2 =-12xw^2 put the letters in abc order
Inner- 3w^2 times 4x =-9w^2
Last- 3w^2 times -3w^2 =12xw^2
your anwser now is
16x-12xw^2-9w^2+12xw^2
you collect like term and waa laa
16x-9w^2-144xw^2 always multiply your like terms or as some people call it "big numbers" do NOTHING to your exponets lots of people mess up beause they do thing to thier exponets! just WATCH OUT!
you got your awnser